The Ashwood Company has a long-term debt ratio of 0.80 and a current ratio of 1.20. Current liabilities are $960, sales are $5,135, profit margin is 9.40 percent, and ROE is 16.90 percent. What is the amount of the firm's net fixed assets?
Answer #1

current ratio =Current assets/Current liabilities

Current assets=(1.2*960)=$1152

Profit margin=net income/sales

Hence net income=(5135*0.094)=$482.69

ROE=net income/equity

Hence equity=($482.69/0.169)=$2856.153846

Long-term debt ratio=Long-term debt/(Long-term debt+equity)

0.8=Long-term debt/(Long-term debt+2856.153846)

0.8Long-term debt+2284.923077=Long-term debt

Long-term debt=2284.923077/(1-0.8)

=$11424.61539

Total debt=Long-term debt+Current liabilities

=$11424.61539+$960

=$12384.61539

Total assets=Total debt+Total equity

Current assets+fixed assets=$12384.61539+$2856.153846

fixed assets=$12384.61539+$2856.153846-$1152

=$14088.77(Approx).